by which factors must the pressure be increased to compress 8L of nitrogen to 2.5L
1
Expert's answer
2020-05-15T10:22:42-0400

P1V1/T1 = P2V2/T2

P1*8/T1 = P2*2.5/T2

if T2 = 2T1 then 8P1 = 2.5P2/2

P2 = 6.4P1

if T2 = 4T1 then 8P1 = 2.5P2/4

P2 = 12.8P1

So we see that if we increase the temperature twice, the pressure will rise twice!

If the temperature (on the kelvin scale) increases by a certain factor, the gas pressure increases by the same factor.